OpenStudy (soulessdeath06):

3x-2y=-6 x+4y=-16 first one to answer CORRECTLY gets medal

OpenStudy (soulessdeath06):

help

OpenStudy (kitty):

I created a step-by-step solution for you at Calorful: https://goo.gl/euf3oN This is a step by step calculator for algebra questions.

OpenStudy (soulessdeath06):

i can veiw it im only allwoed on this site i am blocked on some sites

OpenStudy (soulessdeath06):

*can't

OpenStudy (retireed):

3x-2y=-6 multiply both sides by 2 6x-4y=-12 add this equation to the lower equation +x+4y=-16 ---------- 7x = -28 notice the y terms cancel to out to zero 7x = -28 solve for x then solve to y